Find energy stored in the capacitor CQuestion 46 Consider the following circuit with & = 12 V & = 6 V R, = 32 R2 = 30 R3 = 5 Q R = 10 Rs = 40 R$ = 22 R2 = 322 Rg
Answer #1

Voltage across capacitor Vz+Vy + V6 = Iz (Rg + R4) + R6 Iy z = I + t = 3A But Is = I Ly = IA Ve = 3X(6) + I(2) = 206 Uz tx 4x